OpenStudy (anonymous):

3 is 0.2 percent of what number

OpenStudy (anonymous):

first convert the percent to decimal form 0.2%=.002. There are two ways to do this problem. 1) What number multiplied by 0.2% gives you 3 \[.002x=3\rightarrow x=\frac{3}{.002} \rightarrow x=1500\] 2) you could also set up a proportion and and solve for x, in this case you would leave it in percent form\[\frac{3}{x}=\frac{0.2%}{100%}\rightarrow \frac{3*100%}{.2}=x \rightarrow x=1500\]